OpenStudy (anonymous):

verify (cosBcotB)/(a-sinB) - 1= cscB

OpenStudy (anonymous):

wrtie CotB= cosB/SinB

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

then simply and bring Sin B to denominator , and inside the brackets, and subtract one from fraction and simply, you shall have the answer

OpenStudy (anonymous):

post your work, so its easy to identify if it has errors

OpenStudy (anonymous):

cos2B/(sinB-sin2B) -1 then what

OpenStudy (anonymous):

its shoud be cos^B/(a*Sin B-Sin^2 B) -1

OpenStudy (anonymous):



OpenStudy (anonymous):

i see you missed the 'a'

OpenStudy (anonymous):

did you type the question correctly?

OpenStudy (anonymous):

yes its verify the identities

OpenStudy (anonymous):

(cosBcotB)/(a-sinB) - 1= cscB is what you typed, what happened to "a" in denominator

OpenStudy (anonymous):



OpenStudy (anonymous):

\[= \cos ^{2}B - (\sin B - \sin ^{2}) /(\sin B - Sin^2 B)\]

OpenStudy (anonymous):

that is where i am stuck

OpenStudy (anonymous):

= cos^2 B +sin^2 B-Sin B / (sin B-sin^2 B)

OpenStudy (anonymous):

=1-sin B / (sin B) (1-Sin B)= 1/Sin B = cosec B = : )

OpenStudy (anonymous):

so 1-sinB/(sinB-sin2B)

OpenStudy (anonymous):

happy ?

OpenStudy (anonymous):

hello are you there ?